OpenStudy (anonymous):

What are s the factored form of b²- 4b -32? A. (B-8)(b-4) B. (B-8)(B+4) C. (B+8)(B-4) D. (B+8)(B+4) What is the factored form of 4k²-28+49? A. (2k-7)² B. (2k-14)² C. (2k+7)(2k-7) D. 4(K+7)(k-7)

rishavraj (rishavraj):

clue for first one : -8 + 4 = -4 -8*(4) = -32

rishavraj (rishavraj):

@pinkgymnast there???

OpenStudy (anonymous):

Sorry yah